# Initial Guessing Bias: How Untrained Networks Favor Some Classes

### v3

@@ Abstract
Understanding and controlling biasing effects in neural networks is crucial for ensuring accurate and fair model performance. In the context of classification problems, we provide a theoretical analysis demonstrating that the structure of a deep neural network (DNN) can condition the model to assign all predictions to the same class, even before the beginning of training, and in the absence of explicit biases. We prove that, besides dataset properties, the presence of this phenomenon, which we call Initial Guessing Bias (IGB), is influenced by model choices including dataset preprocessing methods, and architectural decisions, such as activation functions, max-pooling layers, and network depth. Our analysis of IGB provides information for architecture selection and model initialization. We also highlight theoretical consequences, such as the breakdown of node-permutation symmetry, the violation of self-averaging and the non-trivial effects that depth has on the phenomenon.
Machine Learning, ICML, Deep Learning Theory, Theoretical, Bias, Initialization, Classification, Self-Averaging
@@ 1 Introduction
Figure 1: Initial Guessing Bias (IGB). Consider a task where we classify a binary dataset using an untrained network. Does it assign half of the examples to each class, or does it privilege one class? The answer depends on the model design. In the top left, we classify a binary dataset with an untrained network without IGB. This model will generally assign half of the examples to each class (histogram on the top center). In the bottom left, we classify the same dataset using an untrained network with IGB. In this case, most of the guesses will usually go to one of the two classes (histogram on the bottom center). As an example, we take the dog/cat classes (label 0 / label 11) from CIFAR10, and pass them through an untrained CNN with 2 layers, each followed by pooling. The non-IGB model uses $\tanh$ activations and average pooling, the IGB model uses ReLU and Max pooling. We show in the center-right the distribution over different initializations, $p_{f_{0}}\left({x}\right)$, of the fraction $f_{0}$ of times that each model guessed dog (equivalently, $f_{1}=1-f_{0}$ indicates the fraction of images guessed as cat). While for the non-IGB models, $f_{0}$ is most often 50%, with IGB it most often is either 0% or 100%.
@@
In this paper, we study how different design choices in architecture and in data pre-processing influence the predictions of neural networks at initialization. This leads us to the discovery of a previously unexplored phenomenon, which we illustrate in Fig. 1, where the initial predictions made by untrained neural networks are biased. We name this phenomenon Initial Guessing Bias (IGB). IGB challenges naive assumptions about DNNs and informs crucial design decisions. Our study elucidates how, beyond effects that may be induced by the structure of the data itself, the design of the architecture plays a crucial role in shaping the initial predictive bias of the model. Specifically, our work makes the following contributions:
@@
      Informing critical design choices in architecture, initialization, and data standardization.
@@
      Similarly, we show how data preprocessing procedures are crucial in activating and amplifying IGB.
@@
      We provide empirical evidence of how IGB emerges in an even broader range of practical scenarios (real data, CNNs, ResNets, Vision Transformers, etc.), demonstrating IGB’s prevalence.
@@ 2.1 Bias effects
Bias in machine learning models is a multifaceted issue (Mehrabi et al., 2021). While the term ’bias’ often carries negative connotations, particularly when it leads to performance degradation or fairness concerns, it is important to recognize that not all forms of bias are inherently detrimental (Hagendorff and Fabi, 2023; Pot et al., 2021). Controlled and understood bias can be integral to model functionality, shaping the learning process in beneficial ways. The key lies in the ability to regulate these effects to avoid adverse outcomes.
Most research in this domain has predominantly concentrated on biases stemming from dataset characteristics (Barocas and Selbst, 2016), such as class imbalance (Francazi et al., 2023; Ye et al., 2021), or arising due to specific algorithmic choices (Pessach and Shmueli, 2023). These forms of bias, while critical, represent only a part of the broader spectrum. The impact of biases resulting directly from model design decisions—such as network architecture, activation functions, and initialization strategies—remains comparatively unexplored.
Our work contributes to this less-charted territory by examining how certain design choices in DNNs can introduce biases at the very onset of the training process. This perspective not only broadens our understanding of bias in machine learning but also emphasizes the need for a more comprehensive approach to model development, considering the potential effects of every design decision.
@@ 2.2 Properties of DNNs at initialization
Recently, the study of DNNs at the initialization stage has garnered increasing attention, given that the network’s initial state can significantly influence the training process. For example, the initial distribution of the weights can determine an amplification/decay of the signal coming from the input, or even limit the depth to which signals can propagate through random neural networks (Schoenholz et al., 2016; Hanin and Rolnick, 2018; Glorot and Bengio, 2010; Saxe et al., 2013; Orvieto et al., 2021; Noci et al., 2022). The influence of the initial weights distribution is of even greater importance in over-parametrized networks, for which, weights trained with gradient-based optimizers stay close to initialization (Jacot et al., 2018; Du et al., 2019, 2018; Allen-Zhu et al., 2019; Chizat et al., 2019). Furthermore, the initial state of a DNN can significantly impact its generalization performance, as explored in recent studies (Ramasinghe et al., 2023). While the study of the initial state of neural networks has received increasing attention over the past few years, our own study focuses on the initial bias and therefore differs from past work in some important key aspects which we detail in App. B.
3 Preliminary remarks
Figure 2: Left: IGB and Performance Bound: Diagram of the accessible performance range conditioned on the behavior of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ in a balanced binary dataset, in accordance with Eq. (1). Right: Comparison of the trend of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). Particularly, $\gamma(\mathcal{A},\psi(\chi))\in\mathbb{R}^{+}$ (colormap on the right) provides a measure of the level of IGB: the higher the value of $\gamma(\mathcal{A},\psi(\chi))$, the higher the level of IGB (see Sec. 4 for more details). The curves show a consistent pattern with the diagram on the left. They also demonstrate that the time for IGB absorption increases with the level of IGB itself. The simulations were conducted on an MLP-mixer using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting and additional experiments with more architectures are provided in App. I.2.
@@ 3 Preliminary remarks
As we will show in this paper, IGB is a pervasive phenomenon emerging in untrained models. Therefore, a natural question arises: what are the consequences of IGB in terms of training and generalization? Fully addressing this question and determining whether IGB has beneficial or detrimental effects on network training is complex, as the answer might vary with the problem (data, architecture, optimizer, choices of hyperparameter, and so on might all have confounding effects). A comprehensive study of this kind is beyond the scope of this work. However, we can start with a simple observation; a discrepancy between the fractions of guesses, $f_{c}\left(\mathcal{W}_{t}\right)$}, and the actual class proportions in the dataset, sets a limit on the achievable performance. For instance, considering a balanced dataset with $N_{C}$ classes, we find an upper bound on the accuracy (Fig. 2 (left)):
@@ 4 Emergence of IGB: a first insight into the phenomenon
Multi-Layer Perceptron (MLP) submodules are widely integrated into various architectures (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2020); moreover, recent advances, such as the MLP-mixer (Tolstikhin et al., 2021), highlight MLPs’ capability to achieve competitive performance, suggesting their untapped potential, especially in large dataset applications. Our theoretical analysis centers on understanding the intricacies of MLP architectures. Formally, we consider an MLP as a set of $D$ inputs $\{\boldsymbol{\xi}^{(a)}\}_{a=1}^{D}$ that propagate through a set of $L$ hidden layers, until they reach the output layer, composed of a set of two nodes (one for each class), $\left\{O^{(c)}\right\}_{c=0,1}\,$ (see App. A for more details). Each input is classified by selecting the class $c$ with the largest output value.
@@
The choice of random unstructured data as input is common in the literature (Pennington and Worah, 2018; Koehler et al., 2021; Loureiro et al., 2021; Mignacco et al., 2020) as it simplifies theoretical analyses. In our case, this setting has the additional benefit of isolating the analysis from sources of IGB that are potentially embedded in the data structure. To support this assertion, we present empirical results in App. I.1 using real data. These results illustrate that correlations within the data can exacerbate the impact of IGB.
We also note that the Gaussian initialization employed in our analysis is the standard Kaiming initialization (He et al., 2015) but the analysis can be adapted to other initialization schemes . More specifically, our analysis applies to any set of independent weights drawn from a centered distribution with finite variance, with the scaling of the variance being the crucial element.
@@
In our analysis, the random variables (r.v.s) we consider are essentially functions of two independent sources of randomness: the dataset, $\chi$, and the set of initialized weights, $\mathcal{W}$. We denote the cumulative distribution function (c.d.f.) of a r.v. $X$ as $F_{X}\left({x}\right)$ and its probability density function (p.d.f.) as $p_{X}\left({x}\right)$. For variables dependent on both sources of randomness, we specify the active source in the notation. For instance, given $X(\mathcal{W},\chi)$:
$F_{X}^{(\chi)}\left({x}\right)=\mathbb{P}\left(X<x|\mathcal{W}\right)\,,\,\,p_{X}^{(\chi)}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}^{(\chi)}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}}$ (2)
represent the c.d.f. and p.d.f. of $X$ for a fixed set of weights $\mathcal{W}$. As we need to average over both the dataset and the weights, we use the concise notations
@@
Consider a Gaussian distributed dataset processed through an MLP with $L$ hidden layers and weights initialized according to the Kaiming normal scheme (with null bias weights). In the limit of infinite width, the distribution of an output node $O^{(c)}$, at initialization, converge to: $p_{O^{(c)}}^{(\chi)}\left({x}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(x;\mu_{c},\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right),$ (5) where $|\mathcal{W}|$ indicates the cardinality of the set $\mathcal{W}$ and, for compactness, we denoted by $|\mathcal{W}|\rightarrow\infty$ the limit where the number of neurons, $N_{l}$, of each hidden layer $l\in\{0,\dots,L\}$ tends to infinity.
@@
Given an architecture $\mathcal{A}$ and a preprocessed dataset $\psi(\chi)$, we have an absence of IGB if and only if, w.h.p., $\lim_{D\rightarrow\infty}f_{c}\left(\mathcal{W}\right)=\frac{1}{N_{C}},\;\forall c\in\{0,\dots,N_{C}-1\}.$ (7) We instead have IGB if, in the limit of infinite data points ($D\rightarrow\infty$), we observe a disproportion between the values $\{f_{c}\}$ for different classes.
@@
Consistent with the intuition presented in Sec. 4, a possible measure of the level of IGB is given by the ratio of variances:
@@ 5 IGB: emergence and amplification
This analysis not only identifies the architectural elements critical to the emergence of IGB but also quantifies their influence. [^1] We note that, to derive $p_{f_{0}}\left({x}\right)$, we work in the infinite-width limit. This is however not a necessary condition for the appearing of IGB (see e.g. experiments with finite-width models in App. I), but rather an assumption that allows us to derive the theoretical curves.
@@ 5.1 Single hidden layer
    where, in our notation, $\rho^{(m)}\left({\cdot}\right)$ indicates a pooling function with kernel size equals to $m$, $S_{l}^{m}$ indicate the $l^{\mathrm{th}}$ subgroup of $m$ nodes. In particular, for the Max-pooling
@@
Note how, for the output layer, permutation symmetry in nodes corresponds to symmetry between classes. The latter is always preserved at the ensemble level, in fact in all three cases $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$, where $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu\equiv\int p_{\mathcal{W}}\left({\tilde{\mathcal{W}}}\right)f_{0}\left(\tilde{\mathcal{W}}\right)d\tilde{\mathcal{W}}$ indicate the average of $f_{0}\left(\mathcal{W}\right)$ over the ensemble of weight initializations. However, while in the linear case, the symmetry between classes is also conserved on the single element of the ensemble, [^4] In this case, we have $\lim_{D\rightarrow\infty}f_{0}\left(\mathcal{W}\right)=\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$. in the other two cases the single realization diverges, w.h.p. from the mean over the ensemble. NSB thus results in a breaking of the symmetry between classes and a consequent self-averaging breakdown for the observable $f_{0}$. This difference between the two estimates is not due to finite network size effects; our predictions (theoretical curves in Fig. 5) are, in fact, formally exact in the limit of infinite size.
@@
Consider a Gaussian distributed dataset processed through a single hidden layer perceptron, in the limit of infinite width hidden layer, and whose weights are initialized according to the Kaiming normal scheme (with null bias weights). Then, we will have absence of IGB in the network if, and only if, for the generic $i^{\text{th}}$ node of the hidden layer: $\left\langle{\rho^{(1;m)}_{i}}\right\rangle=0\;,\;\;\forall i.$ (16) Or conversely, IGB will emerge if, and only if $\left\langle{\rho^{(1;m)}_{i}}\right\rangle\neq 0\;,\;\;\forall i.$ (17)

### v4

@@ Abstract
Understanding and controlling biasing effects in neural networks is crucial for ensuring accurate and fair model performance. In the context of classification problems, we derive a theoretical analysis demonstrating that the structure of a deep neural network (DNN) can condition the model to assign all predictions to the same class, even before the beginning of training, and in the absence of explicit biases. We prove that, besides dataset properties, the presence of this phenomenon, which we call Initial Guessing Bias (IGB), is influenced by model choices including dataset preprocessing methods, and architectural decisions, such as activation functions, max-pooling layers, and network depth. Our analysis of IGB provides valuable information for selecting network architectures and initializing models. We also highlight theoretical consequences, such as a breakdown of node-permutation symmetry, a violation of self-averaging and non-trivial effects that depth has on IGB.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Initial Guessing Bias (IGB). Consider a task where we classify a binary dataset using an untrained network. Does it assign half of the examples to each class, or does it privilege one class? The answer depends on the model design. In the top-left, we classify a binary dataset with an untrained network without IGB. This model will generally assign half of the examples to each class (histogram on the top-center). In the bottom-left, we classify the same dataset using an untrained network with IGB. In this case, most of the guesses will usually go to one of the two classes (histogram on the bottom-center). As an example, we take the dog/cat classes (label 0 / label 11) from CIFAR10, and pass them through an untrained CNN with 2 layers, each followed by pooling. The non-IGB model uses $\tanh$ activations and average pooling, the IGB model uses ReLU and max pooling. We show in the center-right the distribution over different initializations, $p_{f_{0}}\left({x}\right)$, of the fraction $f_{0}$ of times that each model guessed dog (equivalently, $f_{1}=1-f_{0}$ indicates the fraction of images guessed as cat). While for the non-IGB models, $f_{0}$ is most often 50%, with IGB it most often is either 0% or 100%.
@@
In this paper, we study how different choices in architecture design and data pre-processing influence the predictions of neural networks at initialization. This leads us to the discovery of a previously unexplored phenomenon, which we illustrate in Fig. 1, where the initial predictions made by untrained neural networks are biased. We name this phenomenon Initial Guessing Bias (IGB). IGB challenges naive assumptions about DNNs and informs crucial design decisions. Our study elucidates how, beyond effects that may be induced by the structure of the data itself, the design of the model plays a crucial role in shaping the initial predictive bias of the model. Specifically, our work makes the following contributions:
@@
      Guiding critical design choices in terms of architecture, initialization, and data standardization.
@@
      We show how data preprocessing procedures are crucial in activating and amplifying IGB.
@@
      We provide empirical evidence of the emergence of IGB in a broader range of practical scenarios, including real data, and a wide spectrum of architectures (e.g., CNNs, ResNets, Vision Transformers), demonstrating the prevalence of IGB.
    • $\circ$
      Besides the theoretical setting employing untrained architectures, IGB is also present in the context of transfer learning, where only the head of the model is randomly initialized while the backbone is pre-trained.
@@ 2.1 Bias effects
Bias in machine learning models is a multifaceted issue (Mehrabi et al., 2021). While the term ’bias’ often carries negative connotations, particularly when it leads to performance degradation (Francazi et al., 2023; Engstrom et al., 2020) or fairness concerns (Torralba and Efros, 2011), it is important to recognize that not all forms of bias are inherently detrimental (Hagendorff and Fabi, 2023; Pot et al., 2021). When controlled, biases can be essential to the learning process of a good model. The key lies in the ability to regulate these effects to avoid adverse outcomes.
Most research in this domain has predominantly concentrated on biases stemming from dataset characteristics (Barocas and Selbst, 2016), such as class imbalance (Francazi et al., 2023; Ye et al., 2021; Panigrahi and Zhu, 2024), or arising due to specific algorithmic choices (Pessach and Shmueli, 2023). These forms of bias, while critical, represent only a part of the broader spectrum. The impact of biases resulting directly from model design decisions—such as network architecture, activation functions, and initialization strategies—remains comparatively unexplored.
Our work contributes to this less-charted territory by examining how certain design choices in DNNs can introduce biases at the very onset of the training process. This perspective not only broadens our understanding of bias in machine learning but also emphasizes the need for a more comprehensive approach to model development that considers the potential effects of every design decision.
@@ 2.2 Properties of DNNs at initialization
Recently, the study of DNNs at the initialization stage has garnered increasing attention, given that the network’s initial state can significantly influence the training process. For example, the initial distribution of the weights can determine an amplification/decay of the signal coming from the input, or even limit the depth to which signals can propagate through random neural networks (Schoenholz et al., 2016; Hanin and Rolnick, 2018; Glorot and Bengio, 2010; Saxe et al., 2013; Orvieto et al., 2021; Noci et al., 2022). The initial state of a DNN also significantly impacts generalization performance (Ramasinghe et al., 2023). While the study of the initial state of neural networks has received increasing attention over the past few years, our own study focuses on the initial bias and therefore differs from past work in some important key aspects which we detail in App. B.
3 Preliminaries
Figure 2: Left: IGB and Performance Bound: Diagram of the accessible performance range conditioned on the behavior of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ in a balanced binary dataset, in accordance with Eq. (1). Right: Comparison of the trend of $\max_{c}\{f_{c}\left(\mathcal{W}_{t}\right)\}$ with that of accuracy during the learning dynamics, varying with the level of guessing bias at initialization (IGB). Particularly, $\gamma(\mathcal{A},\psi(\chi))\in\mathbb{R}^{+}$ (colormap on the right) provides a measure of the level of IGB: the higher the value of $\gamma(\mathcal{A},\psi(\chi))$, the higher the level of IGB (see Sec. 4 for more details). The curves show a consistent pattern with the diagram on the left. They also demonstrate that the time for IGB absorption increases with the level of IGB itself. The simulations were conducted on an MLP-mixer using a binary dataset (dogs vs cats from CIFAR) as input; more details on the setting and additional experiments with more architectures/datasets are provided in App. I.2.
Definition IGB
@@ 3 Preliminaries
Beyond initialization
As we will show in this paper, IGB is a pervasive phenomenon emerging in untrained models. Therefore, a natural question arises: what are the consequences of IGB in terms of training and generalization? Fully addressing this question and determining whether IGB has beneficial or detrimental effects on network training is complex as the answer might vary with the problem (data, architecture, optimizer, choices of hyperparameter, and so on might all have confounding effects). A comprehensive study of this kind is beyond the scope of this work. However, we can start with a simple observation; a discrepancy between the fractions of guesses, $f_{c}\left(\mathcal{W}_{t}\right)$}, and the actual class proportions in the dataset, sets a limit on the achievable performance. For instance, considering a balanced dataset with $N_{C}$ classes, we can derive the following upper bound on the accuracy (Fig. 2 (left)):
@@
Methods for controlling IGB
These experiments highlight the importance of thoroughly understanding the phenomenon to formulate effective control strategies. The analysis we present offers several practical methods to regulate—either increase or decrease—IGB based on practitioners’ needs. In particular, the main design choices outlined are:
  • The choice of activation function: Thm. 5.1, and App. F analyse how the choice of the activation function determines the emergence of IGB. This offers practical guidelines on adapting activation functions to either mitigate or induce IGB. Detailed strategies are described in App.F.3 and exemplified in Fig.F.1, showing that simply adding an offset to the ReLU function can significantly reduce IGB.
  • The choice of data standardization: Thm. 5.2 describe how the choice of data pre-processing can be exploited to tune IGB. Beyond the theoretical derivation in App. G.1, we use these results in our experiments to adjust the level of IGB and compare different cases (see Fig. 2 Fig. I.4 and Fig. I.5).
  • The choice of pooling layer: Thm. 5.2, and App. G.2 demonstrate how the max pool, depending on the chosen kernel size, can regulate the level of IGB, with larger kernel sizes amplifying the phenomenon.
  • The depth of the network: Thm. 5.2, and App. G.3.1 illustrate that while increasing the network’s depth does not cause IGB to emerge, it can amplify IGB if it is already present.
Similarly, many other design choices can be analyzed using the framework presented to determine control strategies for the level of IGB. For example, the choice of the temperature of the SoftMax function (see App. E.2.1 for more details).
@@ 4 Emergence of IGB: a first insight into the phenomenon
Multi-Layer Perceptron (MLP) submodules are widely integrated into various architectures (He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2020). Moreover, recent advances, such as the MLP-mixer (Tolstikhin et al., 2021), highlight MLPs’ capability to achieve competitive performance, suggesting their untapped potential, especially in large dataset applications. Our theoretical analysis centers on understanding the intricacies of MLP architectures. Formally, we consider an MLP as a set of $D$ inputs $\{\boldsymbol{\xi}^{(a)}\}_{a=1}^{D}$ that propagate through a set of $L$ hidden layers, until they reach the output layer, composed of a set of two nodes (one for each class), $\left\{O^{(c)}\right\}_{c=0,1}\,$ (see App. A for more details). Each input is classified by selecting the class $c$ with the largest output value.
@@
The choice of random unstructured data as input is common in the literature (Pennington and Worah, 2018; Koehler et al., 2021; Loureiro et al., 2021; Mignacco et al., 2020) as it simplifies theoretical analyses. In our case, this setting has the additional benefit of isolating the analysis from sources of IGB that are potentially embedded in the data structure. To support this assertion, we present empirical results using real data in App. I.1. These results illustrate that correlations within the data can exacerbate the impact of IGB. We follow the common practice of initializing bias parameters to zero, as noted in seminal works (Glorot and Bengio, 2010; He et al., 2015). By considering unstructured data i.i.d. across different classes and DNNs with null biases, we create a highly symmetric setting, making the presence of predictive bias more counter-intuitive. Extending our analysis to include non-null bias parameters is straightforward and discussed in App. G.3.2. Similarly, App. D.4 discusses how the analysis can be extended to include datasets with classes that are not identically distributed.
We also note that the Gaussian initialization employed in our analysis is the standard Kaiming initialization (He et al., 2015) but the analysis can be adapted to other initialization schemes. More specifically, our analysis applies to any set of independent weights $\{w^{(l)}_{ij}\}$ drawn from a centered distribution with variance $\mathcal{O}(1/N_{l})$.
@@
In our analysis, the random variables (r.v.s) we consider are essentially functions of two independent sources of randomness: the dataset, $\chi$, and the set of initialized weights, $\mathcal{W}$. We denote the cumulative distribution function (c.d.f.) of a r.v. $X$ as $F_{X}\left({x}\right)$ and its probability density function (p.d.f.) as $p_{X}\left({x}\right)$. For variables dependent on both sources of randomness, we specify the active source in the notation. For instance, given a function of two independent sets of random variables, $X(\mathcal{W},\chi)$:
$F_{X}^{(\chi)}\left({x}\right)=\mathbb{P}\left(X<x|\mathcal{W}\right)\,,\,\,p_{X}^{(\chi)}\left({x^{\prime}}\right)=\frac{d}{dx}F_{X}^{(\chi)}\left({x}\right)\Bigr|_{\begin{subarray}{c}x=x^{\prime}\end{subarray}},$ (2)
represents the c.d.f. and p.d.f. of $X$ for a fixed set of weights $\mathcal{W}$. As we need to average over both the dataset and the weights, we use the concise notations
@@
Consider a Gaussian distributed dataset processed through an MLP with $L$ hidden layers and weights initialized according to the Kaiming normal scheme (with null bias weights). In the limit of infinite width, the distribution of an output node $O^{(c)}$, at initialization, converges to: $p_{O^{(c)}}^{(\chi)}\left({x}\right)\xrightarrow{|\mathcal{W}|\rightarrow\infty}\mathcal{N}\left(x;\mu_{c},\textrm{Var}_{\chi}\left({O^{(c)}}\right)\right),$ (5) where $|\mathcal{W}|$ indicates the cardinality of the set $\mathcal{W}$ and, for compactness, we denoted by $|\mathcal{W}|\rightarrow\infty$ the limit where the number of neurons, $N_{l}$, of each hidden layer $l\in\{0,\dots,L\}$ tends to infinity.
@@
Given an architecture $\mathcal{A}$ and a preprocessed dataset $\psi(\chi)$, we have an absence of IGB if and only if, w.h.p., $\lim_{D\rightarrow\infty}f_{c}\left(\mathcal{W}\right)=\frac{1}{N_{C}},\;\forall c\in\{0,\dots,N_{C}-1\}.$ (7) We instead have a presence of IGB if, in the limit of infinite data points ($D\rightarrow\infty$), we observe a disproportion between the values $\{f_{c}\}$ for different classes.
@@
Consistently with the intuition presented in Sec. 4, a possible measure of the level of IGB is given by the ratio of variances:
@@ 5 IGB: emergence and amplification
This analysis not only identifies the architectural elements critical to the emergence of IGB but also quantifies their influence. [^1] We note that, to derive $p_{f_{0}}\left({x}\right)$, we work in the infinite-width limit. This is however not a necessary condition for the appearance of IGB (see e.g. experiments with finite-width models in App. I), but rather an assumption that allows us to derive the theoretical curves.
@@ 5.1 Single hidden layer
    where, in our notation, $\rho^{(m)}\left({\cdot}\right)$ indicates a pooling function with kernel size equals to $m$, $S_{l}^{m}$ indicate the $l^{\mathrm{th}}$ subgroup of $m$ nodes. In particular, for the max pooling
@@
Note how, for the output layer, permutation symmetry in nodes corresponds to symmetry between classes. The latter is always preserved at the ensemble level, [^4] Here, ”ensemble” refers to all possible random initializations specified by the chosen scheme (e.g., Kaiming Normal). in fact in all three cases $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$, where $\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu\equiv\int p_{\mathcal{W}}\left({\tilde{\mathcal{W}}}\right)f_{0}\left(\tilde{\mathcal{W}}\right)d\tilde{\mathcal{W}}$ indicate the average of $f_{0}\left(\mathcal{W}\right)$ over the ensemble of weight initializations. However, while in the linear case, the symmetry between classes is also conserved on the single element of the ensemble, [^5] In this case, we have $\lim_{D\rightarrow\infty}f_{0}\left(\mathcal{W}\right)=\mkern 1.5mu\overline{\mkern-1.5muf_{0}\left(\mathcal{W}\right)\mkern-1.5mu}\mkern 1.5mu=1/2$. in the other two cases the single realization diverges, w.h.p. from the mean over the ensemble. NSB thus results in a breaking of the symmetry between classes and a consequent self-averaging breakdown for the observable $f_{0}$. This difference between the two estimates is not due to finite network size effects; our predictions (theoretical curves in Fig. 5) are, in fact, formally exact in the limit of infinite size.
@@
Consider a Gaussian distributed dataset processed through a single hidden layer perceptron, in the limit of infinite width hidden layer, and whose weights are initialized according to the Kaiming normal scheme (with null bias weights). Then, we will have an absence of IGB in the network if, and only if, for the generic $i^{\text{th}}$ node of the hidden layer: $\left\langle{\rho^{(1;m)}_{i}}\right\rangle=0\;,\;\;\forall i.$ (16) Or conversely, IGB will emerge if, and only if $\left\langle{\rho^{(1;m)}_{i}}\right\rangle\neq 0\;,\;\;\forall i.$ (17)